40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 11 03:48:46 2014, Jenne, Update, LSC, 3f DRMI SensMatMeas_10Nov2014_DRMI.pngSensMatMeas_10Nov2014_PRMI.png
    Reply  Tue Nov 11 21:41:09 2014, Koji, Update, LSC, 3f DRMI sensing mat DRMI_radar.pdf
       Reply  Wed Nov 12 03:22:23 2014, Jenne, Update, LSC, 3f DRMI sensing mat 8x
Message ID: 10696     Entry time: Tue Nov 11 03:48:46 2014     Reply to this: 10698
Author: Jenne 
Type: Update 
Category: LSC 
Subject: 3f DRMI 

I was able to lock the DRMI on 3f signals this evening, although the loops are not stable, and I can hear oscillations in the speakers.  I think the big key to making this work was the placement of the SHP-150 high pass filter at the REFL165 PD, and also the installation of Koji's 110 MHz notch filter just before the amplifier, which is before the demod board for REFL165.  These were done to prevent RF signal distortion.

DRMI 3f:   With DRMI locked on 1f (MICH gain = 1, PRCL gain = -0.05, SRCL gain = 2, MICH = 1*REFL55Q, PRCL = 0.1*REFL11I, SRCL = 1*REFL165I), I excited lines, and found the signs and values for 3f matrix elements.  I was using the same gains, but MICH = 0.5*REFL165Q, PRCL = 0.8*REFL33I and SRCL = -0.2*REFL165I.  Part of the problem is likely that I need to include off-diagonal elements in the input matrix to remove PRCL from the SRCL error signal. 

With the DRMI locked on 1f, I took a sensing matrix measurement.  I don't think we believe the W/ct of the photodiode calibration (we need to redo this), but otherwise the sensing matrix measurement should be accurate.  Since the calibration of the PDs isn't for sure, the relative magnitude for signals between PDs shouldn't be taken as gospel, but within a single PD they should be fine for comparison. 

As a side note, we weren't sure about the MICH -> ITMs balancing, so we checked during a MICH-only, and with the locking apparatus we are unable to measure a difference between 1's for both ITMs in the output matrix, and 1 for ITMX and 0.99 for ITMY.  So, we can't measure 1% misbalances in the actuator, but we think we're at least pretty close to driving pure MICH. 

We kind of expect that SRCL should only be present in the 55 and 165 PDs, although we see it strongly in all of the REFL PDs.  Also, PRCL and SRCL are not both lined up in the I-phase.  So, I invite other people to check what they think the sensing matrix looks like. 

  • The excitation lines (and matching notches) were on from 12:14am (
  • Nov 11 2014 08:14:00 UTC / GPS 1099728856) to 12:20am (
  •  
  • Nov 11 2014 08:20:00 UTC / GPS 
  • 1099729216) for 360sec. 
  • MICH was driven with 800 counts at 675.13 Hz, with +1*ITMY, -1*ITMX. 
  • PRCL was driven with 1000 counts at 621.13 Hz with the PRM. 
  • SRCL was driven with 800 counts at 585.13 Hz using the SRM. 

All 3 degrees of freedom have notches at all 3 of those frequencies in the FM10 of the filter banks (and they were all turned on).  During this time, DRMI was locked with 1f signals. 

DRMI sensing matrix:

 SensMatMeas_10Nov2014_DRMI.png

Earlier in the evening, I also took a PRMI sensing matrix, with the PRMI locked on REFL33 I&Q.  Watch out for the different placement of the plots - I couldn't measure AS55 in the DRMI case, since cdsutils.avg freaked out if I asked for more than 14 numbers (#PDs * #dofs) at a time.

SensMatMeas_10Nov2014_PRMI.png

Rana, Koji and I were staring at the DRMI sensing matrix for a little while, and we aren't sure why PRCL and SRCL aren't co-aligned, and why they aren't orthogonal to MICH.  Do we think it's possible to do something to digitally realign them?  Will the solution that we choose be valid for many CARM offsets, or do we have to change things every few steps (which we don't want to do)? 

Things to work on:

* Reanalyze DRMI sensing matrix data from 12:14-12:20am. 

* Make a simulated scan of higher order mode resonances in the arm cavities.  Is it possible that on one or both sides of the CARM resonance we are getting HOM resonances of the sidebands? 

* Figure out how to make DRMI 3f loops stable.

* Try CARM offset reduction with DRMI, and / or PRMI on REFL 165.

ELOG V3.1.3-